Nickel metal, Ni (s), was left in a solution of aluminum (III) ions (Al³⁺), iron (II) ions (Fe2⁺), and copper (II) ions (Cu²⁺). What reaction is most likely to be spontaneous? Write a balanced redox reaction equation for the reaction.


1
Expert's answer
2022-01-18T02:55:01-0500

Question:

Nickel metal, Ni (s), was left in a solution of aluminum (III) ions (Al³⁺), iron (II) ions (Fe2⁺), and copper (II) ions (Cu²⁺). What reaction is most likely to be spontaneous? Write a balanced redox reaction equation for the reaction.


Answer:

The standard electrode potential for the oxidation of nickel is 0.257 V. Therefore, for a reaction to be spontaneous, the reduction potential of the anode must be higher than -0.257 V. The reduction potentials of aluminum (III) ions (Al³⁺), iron (II) ions (Fe2⁺), and copper (II) ions (Cu²⁺) are -1.676 V, -0.44 V and 0.3419 V, respectively. As one can see, only copper (II) ions correspond to the criterion. The balanced redox equation for the reaction is:

Ni(s) + Cu2+(aq) "\rightarrow" Ni2+(aq) + Cu(s), E° = 0.3419- (-0.257) = 0.5989 V